OpenStudy (anonymous):

If given y'=.05y-.04y-1000 how does one arrive at y=100,000-90,000e^(.01t)?

OpenStudy (anonymous):

Here's the context for the entire problem, I'll post it incase it'll help.

OpenStudy (freckles):

Well first you have like terms on the right hand side.

OpenStudy (freckles):

\[y'=.01y-1000 \\ \frac{dy}{dx}=.01y-1000 \\ \frac{dy}{.01y-1000}=dx\]

OpenStudy (freckles):

integrate both sides

OpenStudy (anonymous):

How would I integrate that?

OpenStudy (freckles):

you could use a substitution

OpenStudy (freckles):

but it involves knowing \[\int\limits_{}^{}\frac{du}{u} \text{ equals } \ln|u|+C\]

OpenStudy (anonymous):

But how would that given me y=100,000-90,000e^(.01t)?

OpenStudy (freckles):

How about we go through the process and see.

OpenStudy (freckles):

I think you know the y=e^x and y=ln(x) are inverse functions so it is very likely we will use that

OpenStudy (anonymous):

Well, I got the left side to be ln(.01y-1000)C=a (some random constant?)

OpenStudy (freckles):

\[\frac{1}{.01}\ln|0.01y-1000|=x+C\]

OpenStudy (freckles):

Now solve for y.

OpenStudy (anonymous):

I still don't get it. 100[(e^(.01y)-e^(1000)]=x+c

OpenStudy (anonymous):

Does that step even make sense? :(

OpenStudy (freckles):

\[\ln|0.01y-1000|=\frac{1}{100}(x+c) \\ e^{\ln|0.01y-1000|}=e^{\frac{1}{100}(x+c)} \\ 0.01y-1000=e^{\frac{x}{100}}e^{\frac{c}{100}} \\ 0.01y-1000=e^{\frac{x}{100}}K\]

OpenStudy (anonymous):

Y=100,000+Ke^(.01)?

OpenStudy (freckles):

yep it looks fine.

OpenStudy (anonymous):

but it doesn't match the answer the teacher gave :/

OpenStudy (freckles):

You didn't give an initial condition.

OpenStudy (freckles):

That will allow us to find K

OpenStudy (anonymous):

Oh, yeah, I was meaning to type out the entire problem. I'll do it right now.

OpenStudy (anonymous):

A certain lake contains 10,000 fish. They produce at the rate of 5% per year. They die of natural causes at 4% per year. Fishermen catch 1000 fish per year. a) draw a box diagram for the rate of change in the number of fish (I did that and it matched the answer given). b) Write down the corresponding differential equation for the number of fish in the lake, y(t), and solve it. (I got the proper diff eq. but I couldn't solve it). c) are the fish doomed to extinction? How many fish will there be in the lake after a long time (t=H)? d) At what rate can the fish be caught to keep the population stable?

OpenStudy (freckles):

\[y=100000+k e^\frac{t}{100}\] and base on that above we do we have as the int ital condition?

OpenStudy (freckles):

Look at the first sentence

OpenStudy (anonymous):

at t=0, y=10,000

OpenStudy (freckles):

use that to solve for K

OpenStudy (freckles):

\[10000=100000+ke^{0}\]

OpenStudy (anonymous):

Oh! I get it now! :) Thank you!

OpenStudy (freckles):

So you understood the integration part and the solving for y part?

OpenStudy (anonymous):

I get the solving for y part, I'm not sure if I entirely got the integration part. Like, I get the steps you used, but I'm not sure if I could get to that point by myself tomorrow during the exam. :/ Our professor only spent 5 minutes covering this topic, but it's worth about 20% of tomorrow's test.

OpenStudy (freckles):

\[\int\limits_{}^{}\frac{dy}{0.01y-1000}=\int\limits_{}^{}dx\] You are good up to this point?

OpenStudy (freckles):

Because the first thing I did was just separate variables.

OpenStudy (anonymous):

yes. I got that part.

OpenStudy (anonymous):

I just don't understand where the 1/.01 came in while we were integrating.

OpenStudy (freckles):

So the integral on left hand side we can put it in the form as du/u by letting u=0.01y-1000 du=0.01 dy so du/0.01=dy 100du=dy

OpenStudy (freckles):

\[\int\limits_{}^{}\frac{100 du}{u}=\int\limits_{}^{}1 dx \]

OpenStudy (freckles):

There was a constant multiple in front of y that wasn't 1

OpenStudy (freckles):

\[\int\limits_{}^{}\frac{1}{cx+k} dx=\frac{1}{c} \ln|cx+k|+C \] This is because if we let u=cx+k then du= c dx so du/c=dx

OpenStudy (freckles):

\[\int\limits_{}^{}\frac{1}{cx+k} dx=\int\limits_{}^{}\frac{1}{u} \frac{du}{c}=\frac{1}{c}\int\limits_{}^{}\frac{du}{u}=\frac{1}{c}\ln|u|+K=\frac{1}{c}\ln|cx+k|+K\]

OpenStudy (freckles):

our c here of course being the 0.01 and the k being the -1000

OpenStudy (anonymous):

I'm getting so lost... :/

OpenStudy (freckles):

Ok how about with some numbers

OpenStudy (freckles):

\[\int\limits_{}^{} \frac{1}{5x-3} dx \] do you know how to integrate this?

OpenStudy (freckles):

Maybe the decimals throw you off. I'm not sure yet. But if you can integrate this then that is likely.

OpenStudy (anonymous):

(1/5)ln(5x+3)?

OpenStudy (freckles):

(1/5)ln|5x-3|

OpenStudy (freckles):

So you integrated that without doing a substitution?

OpenStudy (anonymous):

Oh! I see. It's the same concept.

OpenStudy (freckles):

and yes it is

OpenStudy (anonymous):

yeah. It was just like a formula I remembered from last year.

OpenStudy (freckles):

ok you were getting mixed up because of the decimal or the variable or both?

OpenStudy (anonymous):

I think the variables, having dy over it was more confusing.

OpenStudy (freckles):

\[\int\limits_{}^{}\frac{1}{0.01y-1000} dy \text{ less confusing to write this way then ? }\]

OpenStudy (anonymous):

Yep!

OpenStudy (freckles):

I also noticed you had trouble with right hand side

OpenStudy (anonymous):

Yeah, I was confused and then I realized it was the same as 1 dx

OpenStudy (freckles):

maybe because it wasn't that you could see the 1 dx=1 dx

OpenStudy (anonymous):

Yep! :)

OpenStudy (freckles):

It is the same thing that happened to the left hand side

OpenStudy (freckles):

I wrote dy but dy can also be written as 1 dy

OpenStudy (freckles):

Anyways I'm glad you understand the integration part better now. And you said you were okay with the solving for y part?

OpenStudy (anonymous):

Yes. I got how that works.

OpenStudy (anonymous):

And finding K.

OpenStudy (anonymous):

Wait...

OpenStudy (anonymous):

Oh never midn e^(c/100) had to be a constant of some sort and you arbitrarily made that K. :) Got it.

OpenStudy (anonymous):

Thank you so much for your patience in explaining.

OpenStudy (freckles):

No problem. Just make sure you ask your questions no matter how dumb they seem to you. No question is dumb.

OpenStudy (freckles):

Later and good luck